OpenStudy (anonymous):

what is the equation of line perpendicular to 3y=-2x+7

hartnn (hartnn):

can u find the slope of given line ?

OpenStudy (anonymous):

is there no other information ?

OpenStudy (aravindg):

there are infinitely many lines unless u specify a point

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

c=

OpenStudy (anonymous):

oh sorry ...it passes through 3,1

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

the equation of any line perpendicular to 3y=-2x+7 is 3x-2y=k

OpenStudy (anonymous):

now solve for k

OpenStudy (anonymous):

how did you find slope?

OpenStudy (aravindg):

slope is given by m where y=mx+c slope of perpendicular line will be -1/m

OpenStudy (anonymous):

@AravindG thank you now i can do it :)

OpenStudy (aravindg):

:) welcomes!!

OpenStudy (anonymous):

so m= (-2/3)?? right??

OpenStudy (anonymous):

well i guess m= -2/3 and 1/m= -3/2

OpenStudy (aravindg):

m=-2/3

OpenStudy (aravindg):

-1/m =3/2

OpenStudy (aravindg):

y-1=(3/2)(x-3)

OpenStudy (anonymous):

oops thanks again ...now i can do the rest ...thank you guys

OpenStudy (aravindg):

2y-2=3x-9

OpenStudy (aravindg):

:)